# Can Static Black Holes in Massive Gravity Serve as Candidates for Aschenbach-Like Phenomena?

### v1

@@ II.2 Time-like Circular Orbits (TCOs)
$2\mathfrak{L}=-\frac{A}{B}=\varrho,$ (8)
@@
[…]
The concept of angular velocity (as measured by an observer at infinity) and $\beta$ in terms of metric parameters 22 .
[…]
$\beta=-r^{2}\Omega^{2}+f(r),$ (11)
@@
An important aspect to consider in the study of TCOs is the value of the $\beta$ function. In regions where $\beta$ is negative, the energy and angular momentum become imaginary, rendering these areas effectively forbidden for the presence of TCOs. Interested readers seeking a deeper understanding of these relationships, and more examples, are encouraged to refer to 22 ; 23 ; 24 .
III Black holes in Massive Gravity: Advantages and disadvantages
@@
The concept of a massive graviton originated with Markus Fierz and Wolfgang Pauli. They formulated a linear theory (Fierz-Pauli theory) describing a massive spin-2 particle in flat spacetime 25 ; 26 . However, this theory suffered from the van Dam-Veltman-Zakharov (vDVZ) discontinuity 25 ; 27 : predictions didn’t smoothly reduce to General Relativity (GR) as the graviton mass approached zero, conflicting with precise solar system tests 25 ; 28 . Vainshtein proposed a resolution. He demonstrated that nonlinear interactions in massive gravity lead to strong effects (”Vainshtein screening”) near massive sources, dynamically suppressing the helicity-0 mode of the graviton and restoring agreement with General Relativity (GR) in the massless limit, despite a finite graviton mass. This mechanism resolved the van Dam-Veltman-Zakharov (vDVZ) discontinuity by showing that linearized approximations break down in strong-field regimes, such as solar system scales. However, a consistent nonlinear completion of massive gravity—free of the Boulware-Deser ghost instability—was only achieved decades later by de Rham, Gabadadze, and Tolley (dRGT theory, 2010) 29 , which explicitly incorporates Vainshtein screening as a core feature. Constructing a consistent, ghost-free nonlinear massive gravity theory proved challenging for decades due to the emergence of Boulware-Deser ghosts (unphysical negative-energy states) in early formulations. The field was revolutionized in 2010 by Claudia de Rham, Gregory Gabadadze, and Andrew Tolley (dRGT massive gravity). They identified a specific class of nonlinear interaction terms—tuned potential terms constructed from invariants of the dynamical metric and the fixed reference metric—which systematically eliminate the Boulware–Deser ghost to all orders in the Hamiltonian formulation. This provided the first consistent four-dimensional (4D), Lorentz-invariant theory of massive gravity in flat spacetime, free of the instability that had plagued earlier attempts. The theory relies on a reference metric (often Minkowski) and a carefully tuned structure of potential terms to preserve constraints that eliminate the ghost. Subsequent analyses by S.F. Hassan and R.A. Rosen confirmed the absence of the Boulware-Deser ghost in dRGT through a canonical ADM formalism 29 ; 30 .
@@
$\clubsuit$ Bimetric Gravity: A closely related and dynamically richer framework involves two interacting metrics: one dynamical (spacetime) and one fixed reference metric (often Minkowski or de Sitter). Bimetric gravity naturally incorporates massive gravity limits and provides a more flexible setting for finding black hole solutions, including rotating ones (analogues of Kerr) which are much harder to find in pure dRGT 30 ; 31 .
Although the formation of a black hole in the form of Massive gravity still follows the same basic principles of general relativity, but Massive gravity modifies how gravity works. The key differences lie in how the collapse proceeds and the resulting structure of the black hole due to the modified gravitational dynamics, means:
@@ 3. Vainshtein Mechanism:
III.1 Bardeen AdS Black Holes in Massive Gravity
In the presence of massive gravity, the action coupled to nonlinear electrodynamics with negative cosmological constant is given by 32 :
$S$ $=$ $\frac{1}{2}\int d^{4}x\sqrt{-g}\Big{[}R-2\Lambda+m_{g}^{2}\sum_{i}^{4}c_{i}\,\mathcal{U}_{i}(g,f)-\frac{1}{4\pi}\mathcal{L}(F)\Big{]},$ (12)
where $R$,$\Lambda$ , $m_{g}$ and $f$ are scalar curvature, cosmological constant, parameter of massive gravity and symmetric tensor respectively, $c_{i}$ are constants and $\mathcal{U}_{i}$ are polynomials of eigenvalues of the ($4\times 4$) matrix $\mathcal{K}_{\nu}^{\mu}=\sqrt{g^{\mu\alpha}f_{\alpha\nu}}$ which can be written as:
@@ III.1 Bardeen AdS Black Holes in Massive Gravity
${\cal{L}}(F)=\frac{1}{2se^{2}}\left(\frac{\sqrt{2e^{2}F}}{1+\sqrt{2e^{2}F}}\right)^{\frac{5}{2}},$
with $\left[\mathcal{K}\right]=\left[\mathcal{K}_{\mu}^{\mu}\right]$ and $\mathcal{L}(F)$ a function of $F\equiv\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ is the Lagrangian density of Bardeen source 33 .
The spherically-symmetric static metric line element of such a spacetime is given by the following 34 :
$f(r)=1-\frac{2Mr^{2}}{(r^{2}+q^{2})^{\frac{3}{2}}}+\frac{r^{2}}{l^{2}}+m_{g}^{2}\left(C^{2}c_{2}+\frac{Cc_{1}r}{2}\right),$ (13)
where $M$, $l$ and $q$ are mass, AdS radius and magnetic charge.
As stated in the past, since the behavior of black holes in a massive gravitational structure is deeply dependent on parameter values, it is first useful to consider the effect of changes in each parameter on the radius of the event horizon.
@@
The behavioral differences of the metric function, driven by parametric variations, are clearly observable in Fig. (1). As evident from panels Fig. (1) and Fig. (1), the interplay between the black hole mass and the graviton mass significantly influences the structure of the solution. When these two masses approach one another in magnitude, the system tends toward a naked singularity regime, implying a breakdown of the event horizon. A similar trend is observed with increasing electric charge Fig. (1): higher values of charge progressively push the system toward singular behavior, eliminating the possibility of horizon formation. Furthermore, increasing the curvature radius of the spacetime—alternatively interpreted as the AdS radius—directly affects the radius of the event horizon. As depicted in Fig. (1), larger values of $l$ expand the horizon radius; however, excessive growth (e.g., $l=20$) leads to the degeneration of the metric function into a single-root structure, effectively eliminating the outer horizon. The influence of the parameter $c_{1}$ appears negligible within small domains. Nevertheless, beyond a critical threshold, its contribution becomes substantial, causing the Cauchy horizon and the event horizon to approach one another Fig. (1), which signals a reduction in the inter-horizon region.
@@
$M=\frac{\left(q^{2}+r^{2}\right)^{\frac{3}{2}}\left(2C^{2}l^{2}c_{2}m_{g}^{2}+C\,l^{2}rc_{1}m_{g}^{2}+2l^{2}+2r^{2}\right)}{4r^{2}l^{2}}$ (14)
@@
$T=\frac{\left(2C\,r^{3}c_{1}m_{g}^{2}+\left(2C^{2}c_{2}m_{g}^{2}+2\right)r^{2}-C\,q^{2}rc_{1}m_{g}^{2}-4q^{2}\left(C^{2}c_{2}m_{g}^{2}+1\right)\right)l^{2}+6r^{4}}{8r\,l^{2}\left(q^{2}+r^{2}\right)\pi}$ (15)
@@
$\varsigma$ $=$ $\frac{\partial M}{\partial T}\bigg{|}_{r_{h}}=\left(\frac{\partial M}{\partial r}\right)_{r_{h}}\left(\frac{\partial r}{\partial T}\right)_{r_{h}}.$ (16)
@@
- A negative heat capacity indicates instability , as established in the literature 35 ; 36 ; 37 .
@@
By plotting the heat capacity equation with respect to the chosen parametric values Fig. (2), we clearly observe thermodynamically stable behavior for the model within the investigated regime. This analysis reinforce the black hole nature of the configuration and lays the foundation for studying phenomena such as non-monotonic angular velocity profile in the context of Aschenbach-like effects. We now turn to its geodesic behavior of the model. As previously noted, by examining the general form of the metric function as outlined in Eq. (1), one can predict two distinct regimes for characteristic behaviors of TCOs based on the stability of the photon sphere:
@@
Considering the Eq. (2) and Eq. (5) we obtain the effective potential $H$ and components of the vector field $\phi(r,\Theta)$ as:
$H=\frac{\sqrt{1-\frac{2r^{2}M}{\left(q^{2}+r^{2}\right)^{\frac{3}{2}}}+\frac{r^{2}}{l^{2}}+\left(\frac{1}{2}rc_{1}C+C^{2}c_{2}\right)m_{g}^{2}}}{\sin\!\left(\theta\right)r},$ (17)
$\phi^{r_{h}}=-\frac{\left(\left(c_{2}m_{g}^{2}C^{2}+\frac{1}{4}rc_{1}m_{g}^{2}C+1\right)\left(q^{2}+r^{2}\right)^{\frac{5}{2}}-3M\,r^{4}\right)\csc\!\left(\theta\right)}{\left(q^{2}+r^{2}\right)^{\frac{5}{2}}r^{2}},$ (18)
$\phi^{\Theta}=-\frac{\sqrt{\frac{\left(\left(c_{2}m_{g}^{2}C^{2}+\frac{1}{2}rc_{1}m_{g}^{2}C+1\right)l^{2}+r^{2}\right)\left(q^{2}+r^{2}\right)^{\frac{3}{2}}-2r^{2}M\,l^{2}}{\left(q^{2}+r^{2}\right)^{\frac{3}{2}}l^{2}}}\,\cot\!\left(\theta\right)\csc\!\left(\theta\right)}{r^{2}}.$ (19)
Figure 3: Fig (3a): The photon spheres location at $(r,\theta)=(7.7667,1.57)$ and $(r,\theta)=(12.0996,1.57)$ with respect to $M=2,l=1,q=0.5,C=1,c_{1}=-1,c_{2}=1,m_{g}=0.5$ in the $(r-\theta)$ plane of the normal vector field $n$ , (3b): the topological potential H(r) for Bardeen black holes in massive gravity
As illustrated in Fig. (3), while the model indeed exhibits black hole behavior, a notable feature emerges: two distinct photon spheres —an unstable photon sphere (U-PS:corresponding to a maximum of the potential $H$) and a stable photon sphere (S-PS:associated with a local minimum of $H$)—both appear outside the event horizon. Now, if we look at the beta function for this case Fig. (4), we will see that, as mentioned, the region between these two photon spheres has positive beta, and outside it, as stated in the two conditions above, there are forbidden regions due to the negative beta.
Figure 4: Fig (4a): $\beta$ diagram (4b):MSCO localization and space classification for the Bardeen black holes in massive gravity
In Fig. (4), For potential $V$,as defined by Eq. (9) we plot the second derivative of the $V$, to investigate the geodesic behavior of spacetime surrounding the black hole. The profile reveals the presence of two marginally stable circular orbits (MSCOs) :
@@
Having established the theoretical framework, physical prerequisites, and resulting conditions, we now turn our attention to the angular velocity profile and examine its behavior within the scope of this model. In our previous work, we thoroughly explored the mechanism underlying the emergence of this phenomenon in static black hole configurations, highlighting the dependencies of angular velocity on several interconnected quantities:
@@
$\begin{split}&\Omega=\frac{Lf}{E\,r^{2}},\\ &E=\frac{f}{\sqrt{\beta}},\\ &L=-\frac{r^{2}\Omega}{\sqrt{\beta}},\\ &\Omega=\sqrt{\frac{f\!\left(r\right)-\beta}{r^{2}}}.\\ \end{split}$ (20)
@@
III.2 Non-linear charged AdS black hole in massive gravity
If we use the following Lagrangian instead of the Bardeen Lagrangian in Eq. (12),
@@ III.2 Non-linear charged AdS black hole in massive gravity
we will arrive at a nonlinear charged AdS model in massive gravity with metric of the spacetime as follows 38 :
$f(r)$ $=$ $1-\frac{2M}{r}e^{-\frac{k}{2r}}+\frac{r^{2}}{l^{2}}+m_{g}^{2}\left(\frac{Cc_{1}r}{2}+C^{2}c_{2}\right),$ (21)
@@
Given that the methodology and analytical framework have been thoroughly outlined in preceding sections, and considering the recurring structure of the study, we will refrain from repeating detailed procedural explanations in the remainder of this work. To maintain clarity and avoid unnecessary fatigue or confusion, from this point onward, we shall focus exclusively on presenting essential findings and critical observations relevant to each parametric scenario.
@@
The behavioral differences of the metric function, driven by parametric variations, are clearly observable in Fig. (6).It can be clearly seen that while changing $l$ and $c_{1}$ does not have a particular effect in terms of metric and geometric horizon, increasing k causes the model to move towards extremality.
@@
Now with respect to Eq. (2) and Eq. (5) for the effective potential $H$ and components of the vector field $\phi(r,\Theta)$ we have:
$H=\frac{\sqrt{1-\frac{2{\mathrm{e}}^{-\frac{k}{2r}}M}{r}+\frac{r^{2}}{l^{2}}+\left(\frac{rc_{1}}{2}+Cc_{2}\right)Cm_{g}^{2}}}{\sin\!\left(\theta\right)r},$ (22)
$\phi^{r_{h}}=-\frac{\left(\frac{M\left(k-6r\right){\mathrm{e}}^{-\frac{k}{2r}}}{2}+r^{2}\left(c_{2}m_{g}^{2}C^{2}+\frac{1}{4}rc_{1}m_{g}^{2}C+1\right)\right)\csc\!\left(\theta\right)}{r^{4}},$ (23)
$\phi^{\Theta}=-\frac{\sqrt{2}\,\sqrt{\frac{-4{\mathrm{e}}^{-\frac{k}{2r}}M\,l^{2}+2r\left(\left(c_{2}m_{g}^{2}C^{2}+\frac{1}{2}rc_{1}m_{g}^{2}C+1\right)l^{2}+r^{2}\right)}{r\,l^{2}}}\,\cot\!\left(\theta\right)\csc\!\left(\theta\right)}{2r^{2}}.$ (24)
Figure 7: Fig (7a): The photon spheres location at $(r,\theta)=(8.0160,1.57)$ and $(r,\theta)=(9.8766,1.57)$ with respect to $M=2,l=1,k=0.5,C=1,c_{1}=-1,c_{2}=1,m_{g}=0.53$ in the $(r-\theta)$ plane of the normal vector field $n$ , (7b): the topological potential H(r) for Non-linear charged AdS black hole in massive gravity
@@
Figure 8: Fig (4a): $\beta$ diagram, $U-PS(bluedot)=8.0160$ and $S-PS(reddot)=9.8766$ (4b): MSCO localization and space classification for theNon-linear charged AdS black hole in massive gravity
@@
III.3 ModMax-dRGT-like Black hole in massive gravity
If we use the following Lagrangian in Eq. (12),we will arrive at a ModMax-dRGT-like Black hole in massive gravity with metric of the spacetime as follows 39 :
$\mathcal{L}(\mathcal{F})=\mathcal{S}\cosh\gamma-\sqrt{\mathcal{S}^{2}+\mathcal{P}^{2}}\sinh\gamma,$ (25)
$f(r)=1-\dfrac{M}{r}-\frac{\Lambda}{3}r^{2}+\frac{q^{2}e^{-\gamma}}{r^{2}}+m_{g}^{2}C\left(\dfrac{c_{1}r}{2}+c_{2}C\right),$ (26)
where $\mathcal{S}=\frac{\mathcal{F}}{4}$ and $\mathcal{P}=\frac{\widetilde{\mathcal{F}}}{4}$, respectively, are a true scalar, and a pseudoscalar and $\mathcal{F}=F_{\mu\nu}F^{\mu\nu}$ is the Maxwell invariant. Also $\Lambda$, is the cosmological constant, $M$ and $\gamma$ are mass and the dimensionless parameter (which is known as the ModMax’s parameter).

### v3

@@ 2.2 Time-like Circular Orbits (TCOs)
$2\mathfrak{L}=-\frac{A}{B}=\varrho,$ (2.8)
@@
$V_{\mathit{eff}\!\left(\varrho\right)}\!\left(r\right)=\varrho+\frac{A}{B}.$ (2.9)
The concept of angular velocity and $\beta$ in terms of metric parameters [19].
$\Omega=\frac{g_{\mathit{tt}}L}{g_{\varphi\varphi}E},$ (2.10)
$\beta=-g_{tt}-2g_{t\varphi}\Omega-g_{\varphi\varphi}\Omega^{2}=-r^{2}\Omega^{2}+f(r),$ (2.11)
@@
An important aspect to consider in the study of TCOs is the value of the $\beta$ function. In regions where $\beta$ is negative, the energy and angular momentum become imaginary, rendering these areas effectively forbidden for the presence of TCOs. Interested readers seeking a deeper understanding of these relationships, and more examples, are encouraged to refer to [19, 20, 21].
3 Black holes in Massive Gravity: Advantages and disadvantages
@@
The concept of a massive graviton originated with Markus Fierz and Wolfgang Pauli. They formulated a linear theory (Fierz-Pauli theory) describing a massive spin-2 particle in flat spacetime [22, 23]. However, this theory suffered from the van Dam-Veltman-Zakharov (vDVZ) discontinuity [22, 24]: predictions didn’t smoothly reduce to General Relativity (GR) as the graviton mass approached zero, conflicting with precise solar system tests [22, 25]. Vainshtein proposed a resolution. He demonstrated that nonlinear interactions in massive gravity lead to strong effects (”Vainshtein screening”) near massive sources, dynamically suppressing the helicity-0 mode of the graviton and restoring agreement with General Relativity (GR) in the massless limit, despite a finite graviton mass. This mechanism resolved the van Dam-Veltman-Zakharov (vDVZ) discontinuity by showing that linearized approximations break down in strong-field regimes, such as solar system scales. However, a consistent nonlinear completion of massive gravity—free of the Boulware-Deser ghost instability—was only achieved decades later by de Rham, Gabadadze, and Tolley (dRGT theory, 2010) [26], which explicitly incorporates Vainshtein screening as a core feature. Constructing a consistent, ghost-free nonlinear massive gravity theory proved challenging for decades due to the emergence of Boulware-Deser ghosts (unphysical negative-energy states) in early formulations. The field was revolutionized in 2010 by Claudia de Rham, Gregory Gabadadze, and Andrew Tolley (dRGT massive gravity). They identified a specific class of nonlinear interaction terms—tuned potential terms constructed from invariants of the dynamical metric and the fixed reference metric—which systematically eliminate the Boulware–Deser ghost to all orders in the Hamiltonian formulation. This provided the first consistent four-dimensional (4D), Lorentz-invariant theory of massive gravity in flat spacetime, free of the instability that had plagued earlier attempts. The theory relies on a reference metric (often Minkowski) and a carefully tuned structure of potential terms to preserve constraints that eliminate the ghost. Subsequent analyses by S.F. Hassan and R.A. Rosen confirmed the absence of the Boulware-Deser ghost in dRGT through a canonical ADM formalism [26, 27].
@@
$\clubsuit$ Bimetric Gravity: A closely related and dynamically richer framework involves two interacting metrics: one dynamical (spacetime) and one fixed reference metric (often Minkowski or de Sitter). Of course, it should be noted that this fixed reference metric is from the perspective of the massive gravity of dRGT, which seems necessary and one can also be considered it as a dynamic reference. Bimetric gravity naturally incorporates massive gravity limits and provides a more flexible setting for finding black hole solutions, including rotating ones (analogues of Kerr) which are much harder to find in pure dRGT [27, 28].
Although the formation of a black hole in the form of Massive gravity still follows the same basic principles of general relativity, but Massive gravity modifies how gravity works. The key differences in both the collapse process and the resulting black hole structure originate from the modified gravitational dynamics of Massive Gravity. This means:
@@ 3. Vainshtein Mechanism:
3.1 Bardeen AdS Black Holes in Massive Gravity
In the presence of massive gravity, the action coupled to nonlinear electrodynamics with negative cosmological constant is given by [29]:
$S$ $=$ $\frac{1}{2}\int d^{4}x\sqrt{-g}\Big[R-2\Lambda+m_{g}^{2}\sum_{i}^{4}c_{i}\,\mathcal{U}_{i}(g,f)-\frac{1}{4\pi}\mathcal{L}(F)\Big],$ (3.1)
@@ 3.1 Bardeen AdS Black Holes in Massive Gravity
[…]
${\cal{L}}(F)=\frac{1}{2sq^{2}}\left(\frac{\sqrt{2e^{2}F}}{1+\sqrt{2e^{2}F}}\right)^{\frac{5}{2}},$
with $\left[\mathcal{K}\right]=\mathcal{K}_{\mu}^{\mu}$ , $\mathcal{L}(F)$ a function of $F\equiv\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ is the Lagrangian density of Bardeen source and the parameter $s$ is related to the magnetic monopole charge $q$ and mass $M$ by relation $s=q/2M$ [30, 31].The spherically-symmetric static metric line element of such a spacetime is given by the following [31]:
$f(r)=1-\frac{2Mr^{2}}{(r^{2}+q^{2})^{\frac{3}{2}}}+\frac{r^{2}}{l^{2}}+m_{g}^{2}\left(C^{2}c_{2}+\frac{Cc_{1}r}{2}\right),$ (3.2)
where $l$ is AdS radius.
@@
As stated in the past, since the behavior of black holes in a massive gravitational structure is deeply dependent on parameter values, it is first useful to consider the effect of changes in each parameter on the radius of the event horizon. The behavioral differences of the metric function, driven by parametric variations, are clearly observable in Fig. (1). As evident from panels Fig. (1) and Fig. (1), the interplay between the black hole mass and the graviton mass significantly influences the structure of the solution. When these two masses approach one another in magnitude, the system tends toward a naked singularity regime, implying a breakdown of the event horizon. A similar trend is observed with increasing electric charge Fig. (1): higher values of charge progressively push the system toward singular behavior, eliminating the possibility of horizon formation. Furthermore, increasing the curvature radius of the spacetime—alternatively interpreted as the AdS radius—directly affects the radius of the event horizon. As depicted in Fig. (1), larger values of $l$ expand the horizon radius; however, excessive growth (e.g., $l=20$) leads to the degeneration of the metric function into a single-root structure, effectively eliminating the outer horizon. The influence of the parameter $c_{1}$ appears negligible within small domains. Nevertheless, beyond a critical threshold, its contribution becomes substantial, causing the Cauchy horizon and the event horizon to approach one another Fig. (1), which signals a reduction in the inter-horizon region.
@@
$M=\frac{\left(q^{2}+r^{2}\right)^{\frac{3}{2}}\left(2C^{2}l^{2}c_{2}m_{g}^{2}+C\,l^{2}rc_{1}m_{g}^{2}+2l^{2}+2r^{2}\right)}{4r^{2}l^{2}}$ (3.3)
@@
$T=\frac{\left(2C\,r^{3}c_{1}m_{g}^{2}+\left(2C^{2}c_{2}m_{g}^{2}+2\right)r^{2}-C\,q^{2}rc_{1}m_{g}^{2}-4q^{2}\left(C^{2}c_{2}m_{g}^{2}+1\right)\right)l^{2}+6r^{4}}{8r\,l^{2}\left(q^{2}+r^{2}\right)\pi}$ (3.4)
@@
$\varsigma$ $=$ $\frac{\partial M}{\partial T}\bigg|_{r_{h}}=\left(\frac{\partial M}{\partial r}\right)_{r_{h}}\left(\frac{\partial r}{\partial T}\right)_{r_{h}}.$ (3.5)
@@
- A negative heat capacity indicates instability , as established in the literature [32, 33, 34].
@@
By plotting the heat capacity equation with respect to the chosen parametric values Fig. (2), we clearly observe thermodynamically stable behavior for the model within the investigated regime. This analysis reinforce the black hole nature of the configuration and lays the foundation for studying phenomena such as non-monotonic angular velocity profile in the context of Aschenbach-like effects. We now turn to its geodesic behavior of the model. As previously noted, by examining the general form of the metric function as outlined in Eq. (2.1), one can predict two distinct regimes for characteristic behaviors of TCOs based on the stability of the photon sphere:
@@
Considering the Eq. (2.2) and Eq. (2.5) we obtain the effective potential $H$ and components of the vector field $\phi(r,\Theta)$ as:
$H=\frac{\sqrt{1-\frac{2r^{2}M}{\left(q^{2}+r^{2}\right)^{\frac{3}{2}}}+\frac{r^{2}}{l^{2}}+\left(\frac{1}{2}rc_{1}C+C^{2}c_{2}\right)m_{g}^{2}}}{\sin\!\left(\theta\right)r},$ (3.6)
$\phi^{r_{h}}=-\frac{\left(\left(c_{2}m_{g}^{2}C^{2}+\frac{1}{4}rc_{1}m_{g}^{2}C+1\right)\left(q^{2}+r^{2}\right)^{\frac{5}{2}}-3M\,r^{4}\right)\csc\!\left(\theta\right)}{\left(q^{2}+r^{2}\right)^{\frac{5}{2}}r^{2}},$ (3.7)
$\phi^{\Theta}=-\frac{\sqrt{\frac{\left(\left(c_{2}m_{g}^{2}C^{2}+\frac{1}{2}rc_{1}m_{g}^{2}C+1\right)l^{2}+r^{2}\right)\left(q^{2}+r^{2}\right)^{\frac{3}{2}}-2r^{2}M\,l^{2}}{\left(q^{2}+r^{2}\right)^{\frac{3}{2}}l^{2}}}\,\cot\!\left(\theta\right)\csc\!\left(\theta\right)}{r^{2}}.$ (3.8)
Figure 3: (3a): The photon spheres location at $(r,\theta)=(7.7667,1.57)$ and $(r,\theta)=(12.0996,1.57)$ with respect to $M=2,l=1,q=0.5,C=1,c_{1}=-1,c_{2}=1,m_{g}=0.5$ in the $(r-\theta)$ plane of the normal vector field $n$ , (3b): the topological potential H(r) for Bardeen black holes in massive gravity
As illustrated in Fig. (3), while the model indeed exhibits black hole behavior, a notable feature emerges: two distinct photon spheres —an unstable photon sphere (U-PS:corresponding to a maximum of the potential $H$) and a stable photon sphere (S-PS:associated with a local minimum of HH)—both appear outside the event horizon. Now, if we look at the beta function for this case Fig. (4), we will see that, as mentioned, the region between these two photon spheres has positive beta, and outside it, as stated in the two conditions above, there are forbidden regions due to the negative beta.
Figure 4: (4a): $\beta$ diagram (4b):MSCO localization and space classification for the Bardeen black holes in massive gravity
In Fig. (4), For potential $V$,as defined by Eq. (2.9) we plot the second derivative of the $V$, to investigate the geodesic behavior of spacetime surrounding the black hole. The profile reveals the presence of two marginally stable circular orbits (MSCOs) :
@@
Having established the theoretical framework, physical prerequisites, and resulting conditions, we now turn our attention to the angular velocity profile and examine its behavior within the scope of this model. Wei in his valuable work [15] and we in our previous work [11], thoroughly explored the mechanism underlying the emergence of this phenomenon in static black hole configurations, highlighting the dependencies of angular velocity on several interconnected quantities:
@@
[…]
@@
3.2 Non-linear charged AdS black hole in massive gravity
If we use the following Lagrangian instead of the Bardeen Lagrangian in Eq. (3.1),
@@ 3.2 Non-linear charged AdS black hole in massive gravity
we will arrive at a nonlinear charged AdS model in massive gravity with metric of the spacetime as follows [35]:
$f(r)$ $=$ $1-\frac{2M}{r}e^{-\frac{k}{2r}}+\frac{r^{2}}{l^{2}}+m_{g}^{2}\left(\frac{Cc_{1}r}{2}+C^{2}c_{2}\right),$ (3.10)
@@
Given that the methodology and analytical framework have been thoroughly outlined in preceding sections, and considering the recurring structure of the study, we will refrain from repeating detailed procedural explanations in the remainder of this work. To maintain clarity and avoid unnecessary fatigue or confusion, from this point onward, we shall focus exclusively on presenting essential findings and critical observations relevant to each parametric scenario. The behavioral differences of the metric function, driven by parametric variations, are clearly observable in Fig. (6). It can be clearly seen that while changing $l$ and $c_{1}$ does not have a particular effect in terms of metric and geometric horizon, increasing k causes the model to move towards extremality.
@@
Now with respect to Eq. (2.2) and Eq. (2.5) for the effective potential $H$ and components of the vector field $\phi(r,\Theta)$ we have:
$H=\frac{\sqrt{1-\frac{2{\mathrm{e}}^{-\frac{k}{2r}}M}{r}+\frac{r^{2}}{l^{2}}+\left(\frac{rc_{1}}{2}+Cc_{2}\right)Cm_{g}^{2}}}{\sin\!\left(\theta\right)r},$ (3.11)
$\phi^{r_{h}}=-\frac{\left(\frac{M\left(k-6r\right){\mathrm{e}}^{-\frac{k}{2r}}}{2}+r^{2}\left(c_{2}m_{g}^{2}C^{2}+\frac{1}{4}rc_{1}m_{g}^{2}C+1\right)\right)\csc\!\left(\theta\right)}{r^{4}},$ (3.12)
$\phi^{\Theta}=-\frac{\sqrt{2}\,\sqrt{\frac{-4{\mathrm{e}}^{-\frac{k}{2r}}M\,l^{2}+2r\left(\left(c_{2}m_{g}^{2}C^{2}+\frac{1}{2}rc_{1}m_{g}^{2}C+1\right)l^{2}+r^{2}\right)}{r\,l^{2}}}\,\cot\!\left(\theta\right)\csc\!\left(\theta\right)}{2r^{2}}.$ (3.13)
Figure 7: (7a): The photon spheres location at $(r,\theta)=(8.0160,1.57)$ and $(r,\theta)=(9.8766,1.57)$ with respect to $M=2,l=1,k=0.5,C=1,c_{1}=-1,c_{2}=1,m_{g}=0.53$ in the $(r-\theta)$ plane of the normal vector field $n$ , (7b): the topological potential H(r) for Non-linear charged AdS black hole in massive gravity
@@
Figure 8: (8a): $\beta$ diagram, UPS(bluedot)=8.0160U-PS(bluedot)=8.0160 and SPS(reddot)=9.8766S-PS(reddot)=9.8766 (8b): MSCO localization and space classification for the Non-linear charged AdS black hole in massive gravity
@@
3.3 ModMax-dRGT-like Black hole in massive gravity
If we use the following Lagrangian in Eq. (3.1),we will arrive at a ModMax-dRGT-like Black hole in massive gravity with metric of the spacetime as follows [36]:
$\mathcal{L}(\mathcal{F})=\mathcal{S}\cosh\gamma-\sqrt{\mathcal{S}^{2}+\mathcal{P}^{2}}\sinh\gamma,$ (3.14)
$f(r)=1-\dfrac{M}{r}-\frac{\Lambda}{3}r^{2}+\frac{q^{2}e^{-\gamma}}{r^{2}}+m_{g}^{2}C\left(\dfrac{c_{1}r}{2}+c_{2}C\right),$ (3.15)
where $\mathcal{S}=\frac{\mathcal{F}}{4}$ and $\mathcal{P}=\frac{\widetilde{\mathcal{F}}}{4}$, respectively, are a true scalar, and a pseudoscalar and $\mathcal{F}=F_{\mu\nu}F^{\mu\nu}$ is the Maxwell invariant. Also $\Lambda$, is the cosmological constant, $M$ and $\gamma$ are mass and the dimensionless parameter (which is known as the ModMax’s parameter).As can be seen in Fig. (10), since in this model the $\gamma$ parameter will have a significant effect on the gravitational behavior of the model, in such a way that small values (e.g. $\gamma<0.6686$) will push the model towards extremalization and larger values will lead to black hole behavior, we therefore plot the metric function in two cases.